Instructor is to examine a population of 20 students to check for homework. If 15 have done their homework and if sample of 4 students is to be randomly chosen, what is the probability that exactly 3 will have done their homework? What is the probability that exactly 3 will have done their homework



1
Expert's answer
2021-10-04T16:21:02-0400

Let x be the number of students who have done their homework

Given Population(N)= 20 ; Sample(n)=4 ; done homework (m)= m

Then the proof of x is

"P(x=k)= \frac{\begin{pmatrix}
 m \\
 k
\end{pmatrix} \begin{pmatrix}
 N-m \\
 n-k
\end{pmatrix}}{\begin{pmatrix}
 N \\
 n
\end{pmatrix}}= \frac{\begin{pmatrix}
 15 \\
 k
\end{pmatrix} \begin{pmatrix}
 20-m \\
 4-k
\end{pmatrix}}{\begin{pmatrix}
 20 \\
 4
\end{pmatrix}}"

P( exactly 3 will have done their work)

"= P(x=3)\\
= \frac{\begin{pmatrix}
 15 \\
 3
\end{pmatrix} \begin{pmatrix}
 20-15 \\
 4-3
\end{pmatrix}}{\begin{pmatrix}
 20 \\
 4
\end{pmatrix}}\\
= \frac{(\frac{15!}{(15-3)! 3!}) (\frac{5!}{(5-1)! 1!})}{(\frac{20!}{(20-4)! 4!})}\\
=\frac{455*5}{4845}\\
=0.47"